Young’s Experiment is performed with wavelength 420nm and the slits are 2mm apart. How far away should you place a screen (in m) so that the closest fringes are 2mm apart?
Answer #1

ya Si Screen Given Y= 2mm L-? wavelength λ= 420nm and silts d= 2mm We know that Y= L/d Hence L= y*d/ λ = (2*2*10-6)/420*10^-9